OpenStudy (safyousuff97):

Task 4—Solving Rational Equations Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. x plus a over ax = b over x Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means.

OpenStudy (wanell):

is that the only 2 parts you have

OpenStudy (safyousuff97):

yea

OpenStudy (wanell):

We have to cross multiply to get: x(x+1) = 2x x^2 + x = 2x x^2 -x = 0 x(x-1)=0 x=0 or x=1 The x=0 solution is extraneous because we are not allowed to divide by 0 (which is what happens in the original equation if x is 0). Do This Look Right Are Not

OpenStudy (safyousuff97):

thank you :)